Variable Seperable


Cos² y dx + Sin² x dy = 0; y (π/4) =(π/4)


1
Expert's answer
2021-08-31T17:17:59-0400

Let us solve the differential equation "\cos^2y\ dx + \sin^2 x\ dy = 0" which is equivalen to "\sin^2 x\ dy = -\cos^2y\ dx."

If "\sin x=0," then "x=\pi n,\ n\in\Z," are the solutions of the differential equation "\cos^2y\ dx + \sin^2 x\ dy = 0". Also if "\cos y=0," then "y=\frac{\pi}{2}+\pi n,\ n\in\Z," are the solutions of the differential equation "\cos^2y\ dx + \sin^2 x\ dy = 0". Since "y(\frac{\pi}{4})\ne\frac{\pi}{4}," we conclude that they are not the solutions of "\cos^2y\ dx + \sin^2 x\ dy = 0, \ y(\frac{\pi}{4})=\frac{\pi}{4}."


Let us divide both parts of this equation by "\sin^2x\cos^2y." Then we get the following differential equation:

"\frac{dy}{\cos^2 y}=-\frac{dx}{\sin^2 x}"


Then

"\int\frac{dy}{\cos^2 y}=-\int\frac{dx}{\sin^2 x}" and hence


"\tan y=\ctg x+C."


Since "y(\frac{\pi}{4})=\frac{\pi}{4}," we have that "\tan \frac{\pi}{4}=\ctg \frac{\pi}{4}+C." It follows that "1=1+C," and hence "C=0."

We conclude that "\tan y=\ctg x=\tan(\frac{\pi}{2}-x)." It follows that "y=\frac{\pi}{2}-x+\pi n,\ n\in\Z," is the family of solutions of the last differential equation. Taking into account that "y(\frac{\pi}{4})=\frac{\pi}{4}," we conclude that "y=\frac{\pi}{2}-x" is the solution of the differential equation "\cos^2y\ dx + \sin^2 x\ dy = 0, \ y(\frac{\pi}{4})=\frac{\pi}{4}."